# Theory and Experiment

### v1

@@ Abstract
Quantum state tomography (QST) reconstructs the density matrix of an unknown quantum state by measuring many identical copies Paris and Rehacek (2004); Häffner et al. (2005). This technique is essential in quantum computing DiVincenzo (1995); O’brien (2007), quantum communication Bennett et al. (1993); Bouwmeester et al. (1997), and quantum metrology Giovannetti et al. (2006, 2011). For a $d$-dimensional state, the standard approach uses a number of measurement bases of order $d^{2}$ and reconstructs the density matrix from the expectation values of measurement operators James et al. (2001); Thew et al. (2002); Hioe and Eberly (1981); Bertlmann and Krammer (2008); Kimura (2003); Mendaš (2006). Each measurement provides both an expectation value and a probability distribution, but existing methods do not fully use all the information available. Making fuller use of the probability data can reduce the number of measurement bases required. Based on this idea, recent studies have shown that only $12d-$ 1 projective measurement bases are sufficient for complete state reconstruction in any $d$-dimensional quantum system Newton and Young (1968); Perlin et al. (2021); Wang et al. (2024).
Another prominent strategy for QST is based on the construction of mutually unbiased bases (MUBs) Ivonovic (1981); Wootters and Fields (1989); Zauner (2011); Durt et al. (2010); Horodecki et al. (2022). In a $d$-dimensional Hilbert space, at most $d+1$ MUBs can exist in theory. If a complete set of $d+1$ MUBs can be constructed for a given dimension, QST can be performed using only these bases. However, current research shows that a complete set of $d+1$ MUBs can be constructed only when $d$ is a power of a prime Bandyopadhyay et al. (2002); Klappenecker and Rötteler (2004); Archer (2005); Brierley et al. (2009); Adamson and Steinberg (2010); Lima et al. (2011); Giovannini et al. (2013). In other dimensions, this construction remains an open problem; for example, for $d=6$, only three MUBs have been found to date Bengtsson et al. (2007); Grassl (2004); Jaming et al. (2009); Butterley and Hall (2007); Brierley and Weigert (2009); Raynal et al. (2011); Goyeneche (2013). As a result, MUB-based QST lacks universality and cannot be implemented in all dimensions. This raises a natural question: for any dimension $d$, can QST be achieved with fewer than $12d-$ 1 measurement bases?
@@
is Hermitian, normalized as $\mathrm{Tr}(\rho)=1$, and contains $d^{2}-1$ independent real parameters. Measuring the state in each orthonormal basis yields $d$ probability values, but at most $d-1$ of them are independent. Therefore, to fully determine the density matrix, at least $d+1$ distinct measurement bases are required.
@@
where $\omega=e^{2\pi i/d}$. The remaining $d-1$ measurement bases are generated by applying $d-1$ different diagonal unitary operators,
@@
where $j=0,1,\ldots,d-1$ and $k=0,1,\ldots,d-2$, with $k=d-1$ excluded since at most $d-1$ probabilities are independent for each measurement basis. Collecting all $q_{k}^{(j)}$ and $g_{a,b}$ ($a\neq b$) into $d(d-1)$-dimensional column vectors $\bm{q}$ and $\bm{g}$, Eq. (9) can be written compactly as
@@
for $t\neq t’$ ’, $t,t’=0,1,\ldots,d-1$, $c=1,2,\ldots,d-1$, $n=0,\pm 1,\pm 2,\ldots$, and $j,m=0,1,\ldots,d-1$ (see Supplemental Material). The off-diagonal elements, collected in $\bm{g}$, are then uniquely determined by
@@
To experimentally verify this QST scheme, we developed a programmable quantum optical platform based on a silicon photonic chip. As shown in Fig. 1, this platform integrates all essential components, from single-photon generation to quantum state measurement. (1) Single-photon generation. A pulsed laser operating at 80 MHz with a central wavelength of 1550.3 nm is injected into the silicon waveguide spiral on the chip, where photon pairs are produced through the process of four-wave mixing. The pump laser is first removed by an asymmetric Mach-Zehnder interferometer (AMZI). Subsequently, a second AMZI separates the generated photons, with the idler photon (wavelength 1553.5 nm) used as a trigger and the signal photon (wavelength 1547.1 nm) routed to the downstream quantum circuits for further processing. (2) Preparation of six-dimensional quantum states. The signal photon then passes through an interferometric network composed of six Mach-Zehnder interferometers (MZIs), where six heaters allow arbitrary amplitude modulation for each path. Each path is also equipped with an independent heater for phase control, enabling the preparation of any pure six-dimensional quantum state. By preparing up to six such pure states and mixing them with appropriate probabilities, any six-dimensional mixed state can be realized. (3) Quantum state measurement and data acquisition. The six-dimensional quantum state then passes through a reconfigurable six-dimensional linear optical network on the chip, consisting of multiple MZIs and heaters. By adjusting the heaters, arbitrary six-dimensional unitary transformations can be realized Clements et al. (2016); Li et al. (2023). In our experiment, seven distinct unitaries are implemented using a variational approach Peruzzo et al. (2014); Xue et al. (2022), corresponding to seven projective measurement bases. The first measurement basis is the computational basis, and the other six are the Fourier basis and its variants under diagonal unitary transformations, where the diagonal phases are set to $\theta_{m}^{(j)}=0.5415\,jm^{2}$, with $j=1,2,\ldots,5$ and $m=0,1,\ldots,5$ aut . After the signal photon passes through the unitary transformation network, both the signal and idler photons are coupled out of the chip via a fiber array and detected by superconducting nanowire single-photon detectors (SNSPDs), from which the probability distributions for each measurement basis are extracted. Using these distributions, the full density matrix of any six-dimensional quantum state can be reconstructed by applying the theoretical scheme described above together with maximum likelihood estimation James et al. (2001); Thew et al. (2002).
@@
Using our scheme, we experimentally reconstructed the density matrices of all three states, as shown in Fig. 2. The fidelities between the reconstructed density matrices $\rho’$ ’ and the ideal states $\rho$ are calculated as
@@
yielding values of 0.9654Âą0.0067plus-or-minus0.96540.00670.9654$\pm$ 0.00670.9654 Âą 0.0067, 0.9698Âą0.0042plus-or-minus0.96980.00420.9698$\pm$ 0.00420.9698 Âą 0.0042, and 0.9761Âą0.0033plus-or-minus0.97610.00330.9761$\pm$ 0.00330.9761 Âą 0.0033, respectively.

### v2

@@ Abstract
Quantum state tomography (QST) reconstructs the density matrix of an unknown quantum state by measuring many identical copies Paris and Rehacek (2004); Häffner et al. (2005). This technique is essential in quantum computing DiVincenzo (1995); O’brien (2007), quantum communication Bennett et al. (1993); Bouwmeester et al. (1997), and quantum metrology Giovannetti et al. (2006, 2011). For a $d$-dimensional state, the standard approach uses a number of measurement bases of order $d^{2}$ and reconstructs the density matrix from the expectation values of measurement operators James et al. (2001); Thew et al. (2002); Hioe and Eberly (1981); Bertlmann and Krammer (2008); Kimura (2003); Mendaš (2006). Each measurement provides both an expectation value and a probability distribution, but existing methods do not fully use all the information available. Making fuller use of the probability data can reduce the number of measurement bases required. Based on this idea, recent studies have shown that only 2d12d-1 projective measurement bases are sufficient for complete state reconstruction in any $d$-dimensional quantum system Newton and Young (1968); Perlin et al. (2021); Wang et al. (2024).
Another prominent strategy for QST is based on the construction of mutually unbiased bases (MUBs) Ivonovic (1981); Wootters and Fields (1989); Zauner (2011); Durt et al. (2010); Horodecki et al. (2022); McNulty and Weigert (2024). In a $d$-dimensional Hilbert space, at most $d+1$ MUBs can exist in theory. If a complete set of $d+1$ MUBs can be constructed for a given dimension, QST can be performed using only these bases. However, current research shows that a complete set of $d+1$ MUBs can be constructed only when $d$ is a power of a prime Bandyopadhyay et al. (2002); Klappenecker and Rötteler (2004); Archer (2005); Brierley et al. (2009); Adamson and Steinberg (2010); Lima et al. (2011); Giovannini et al. (2013). In other dimensions, this construction remains an open problem; for example, for $d=6$, only three MUBs have been found to date Bengtsson et al. (2007); Grassl (2004); Jaming et al. (2009); Butterley and Hall (2007); Brierley and Weigert (2009); Raynal et al. (2011); Goyeneche (2013). As a result, MUB-based QST lacks universality and cannot be implemented in all dimensions. This raises a natural question: for any dimension $d$, can QST be achieved with fewer than 2d12d-1 measurement bases?
@@
is Hermitian, normalized as $\mathrm{Tr}(\rho)=1$, and contains $d^{2}-1$ independent real parameters. Measuring the state in each orthonormal basis yields $d$ probability values, but at most d1d-1 of them are independent. Therefore, to fully determine the density matrix, at least $d+1$ distinct measurement bases are required.
@@
where $\omega=e^{2\pi i/d}$. The remaining d1d-1 measurement bases are generated by applying d1d-1 different diagonal unitary operators,
@@
where $j=0,1,\ldots,d-1$ and $k=0,1,\ldots,d-2$, with k=d1k=d-1 excluded since at most d1d-1 probabilities are independent for each measurement basis. Collecting all $q_{k}^{(j)}$ and $g_{a,b}$ ($a\neq b$) into d(d1)d(d-1)-dimensional column vectors $\bm{q}$ and $\bm{g}$, Eq. (9) can be written compactly as
@@
for $\neq t’, t,t$$t’=0,1,\ldots,d-1,$ $c=1,2,\ldots,d-1$, $n=0,\pm 1,\pm 2,\ldots$, and $j,m=0,1,\ldots,d-1$ (see Supplemental Material). The off-diagonal elements, collected in $\bm{g}$, are then uniquely determined by
@@
To experimentally verify this QST scheme, we developed a programmable quantum optical platform based on a silicon photonic chip. As shown in Fig. 1, this platform integrates all essential components, from single-photon generation to quantum state measurement. (1) Single-photon generation. A pulsed laser operating at 80 MHz with a central wavelength of 1550.3 nm is injected into the silicon waveguide spiral on the chip, where photon pairs are produced through the process of four-wave mixing. The pump laser is first removed by an asymmetric Mach-Zehnder interferometer (AMZI). Subsequently, a second AMZI separates the generated photons, with the idler photon (wavelength 1553.5 nm) used as a trigger and the signal photon (wavelength 1547.1 nm) routed to the downstream quantum circuits for further processing. (2) Preparation of six-dimensional quantum states. The signal photon then passes through an interferometric network composed of six Mach-Zehnder interferometers (MZIs), where six heaters allow arbitrary amplitude modulation for each path. Each path is also equipped with an independent heater for phase control, enabling the preparation of any pure six-dimensional quantum state. By preparing up to six such pure states and mixing them with appropriate probabilities, any six-dimensional mixed state can be realized. (3) Quantum state measurement and data acquisition. The six-dimensional quantum state then passes through a reconfigurable six-dimensional linear optical network on the chip, consisting of multiple MZIs and heaters. By adjusting the heaters, arbitrary six-dimensional unitary transformations can be realized Clements et al. (2016); Li et al. (2023). In our experiment, seven distinct unitaries are implemented using a variational approach Peruzzo et al. (2014); Xue et al. (2022), corresponding to seven projective measurement bases. The first measurement basis is the computational basis, and the other six are the Fourier basis and its variants under diagonal unitary transformations, where the diagonal phases are set to $\theta_{m}^{(j)}=0.5671\,jm^{2}$, with $j=1,2,\ldots,5$ and $m=0,1,\ldots,5$ aut . After the signal photon passes through the unitary transformation network, both the signal and idler photons are coupled out of the chip via a fiber array and detected by superconducting nanowire single-photon detectors (SNSPDs), from which the probability distributions for each measurement basis are extracted. Using these distributions, the full density matrix of any six-dimensional quantum state can be reconstructed by applying the theoretical scheme described above together with maximum likelihood estimation James et al. (2001); Thew et al. (2002).
@@
Using our scheme, we experimentally reconstructed the density matrices of all three states, as shown in Fig. 2. The fidelities between the reconstructed density matrices $\rho$’ and the ideal states $\rho$ are calculated as
@@
yielding values of $0.9654\pm 0.0067$, $0.9698\pm 0.0042$, and $0.9761\pm 0.0033$, respectively.